### v1

# Quantum Algorithms for Matrix Operations Based on Unitary Transformations and Ancillary State Measurements

Abstract
Matrix operations are of great significance in quantum computing, which manipulate quantum states in information processing. This paper presents quantum algorithms for several important matrix operations. By leveraging multi-qubit Toffoli gates and basic single-qubit operations, these algorithms efficiently carry out matrix row addition, row swapping, trace calculation and transpose. By using the ancillary measurement techniques to eliminate redundant information, these algorithms achieve streamlined and efficient computations, and demonstrate excellent performance with the running time increasing logarithmically as the matrix dimension grows, ensuring scalability. The success probability depends on the matrix dimensions for the trace calculation, and on the matrix elements for row addition. Interestingly, the success probability is a constant for matrix row swapping and transpose, highlighting the reliability and efficiency.
Keywords: quantum algorithm, quantum circuit, matrix operation, row addition, row swapping, quantum measurement
1 Introduction
Quantum computations offer exponential speedups over their classical counterparts by leveraging the principles of quantum mechanics. In 1980 Benioff and Manin BP proposed the concept of quantum computing, and Feynman MYI introduced the theoretical framework of a quantum computer in 1982. In 1985 Deutsch Deu proposed the quantum parallel algorithm, showcasing the significant computational advantages of quantum computing over classical methods.
In 1994 Shor Sho1 ; Sho2 presented a quantum algorithm for prime factorization, drawing widespread attention to the advantages of quantum computing in solving certain problems. Grover Gro followed in 1996 with a quantum search algorithm, further demonstrating quantum computing’s potential in data search applications. In 2009 the HHL algorithm HHL ; HHL1 was developed for solving linear systems of equations, greatly advancing the quantum algorithm researches. Several renowned quantum operations, such as quantum Fourier transform QFT1 ; QFT2 ; QFT3 and quantum phase estimation QFT3 ; Wang , have significantly contributed to the field.
Matrix operations are indispensable tools in quantum algorithms, which play key roles in quantum state manipulations in quantum system evolutions, quantum information processing and quantum computing. The design of quantum algorithms and the implementation of quantum circuits often rely heavily on matrix operations. Both Shor’s algorithm and Grover’s search algorithm optimize their computational processes through matrix operations. The HHL algorithm also relies fundamentally on matrix operations and has found numerous applications in matrix-based protocols Wang ; Tak ; Berry , providing new directions and momentum for the study of matrix operations and inspiring the development of more quantum algorithms for matrix computations. Given the broad and significant applications of matrix operations across disciplines, using quantum computing to address algebraic problems related to matrices has become a prominent research focus.
In recent years, researchers have proposed several quantum algorithms for matrix operations. For example, the Trotter approximation formula THF has been utilized to simulate quantum circuits, addressing complex operations such as matrix addition, multiplication and Kronecker products. Quantum algorithms for computing vector inner products ZhaoZ ; SZ_2019 ; QZKW_arxive2022 ; ZQKW_arxiv2022; ZQKW_arxiv2023 ; ZQKW_arxiv2024 have also been developed. These algorithms leverage the superposition and entanglement of qubits and have potential applications in fields like machine learning and data processing. Based on a “sender-receiver” quantum computing model, in 2022 Wen-Tao Qi et al. QZKW_arxive2022 proposed a class of quantum algorithms for matrix operations, including matrix-vector multiplication, matrix-matrix multiplication, matrix addition, determinant computation and matrix inversion. In 2023 A.I. Zenchuk et al. ZQKW_arxiv2023 further developed the idea of using specific types of unitary transformations to implement linear algebra protocols according to multi-qubit Toffoli gates and the simplest single-qubit operations. Auxiliary measurements were employed to remove redundant information and extract desired results. Building on this work, in 2024 A.I. Zenchuk et al. ZQKW_arxiv2024 introduced a method to encode the $N$ rows of a matrix into pure states of $N$ independent quantum subsystems. By selecting appropriate unitary operators, the determinant can be computed efficiently.
However, these quantum algorithms for basic matrix operations still face challenges such as high computational complexity and low success probabilities, which limit their broader applicability.
The elementary row transformations of matrices play significant roles in solving linear systems of equations, determining matrix rank, finding invertible matrices and simplifying determinant computations, with wide-ranging value in both theoretical research and practical applications. In quantum computing, elementary row transformations can simplify matrices for efficient linear equation solving, provide theoretical support for matrix decomposition, advance large-scale matrix operations, and help to interpret the geometric meanings of linear transformations. These contributions offer new insights into the designs of quantum algorithms. Currently, there are no specialized quantum algorithms specifically designed for elementary row transformations of matrices.
Drawing on the foundational ideas from ZQKW_arxiv2023 , this paper proposes algorithms for two elementary row transformations of matrices, transpose and trace of matrices, along with their corresponding quantum circuits. These algorithms collectively involve the following key steps:
1. Construction of the initial state: matrix elements are first encoded into the probability amplitudes of quantum pure states by using tensor product operations.
2. Information separation and labeling: combining multi-qubit control operators with ancillary states, where the ancillary states are used to accurately label the useful and redundant parts of the information, achieving effective separation of the data.
3. State swaping operations: depending on the specific algorithms, C-SWAP operations are flexibly employed to precisely exchange the states of two quantum bits.
4. State superposition processing: Hadamard operators are applied to induce superposition states among quantum bits.
5. Measurement on the ancillary states: auxiliary measurement techniques are used to eliminate redundant information, ensuring that the desired results are accurately reflected in the probability amplitudes of the final state.
The remainder of this paper is organized as follows: Section 2 introduces quantum algorithms for two types of elementary matrix transformations. Section 3 presents quantum algorithms for computing the trace and the transposition of a matrix. Conclusion is given in section 4.
2 Two types of elementary row transformations
This section presents quantum algorithms for two types of elementary row transformations of an $N\times M$ matrix $A=\{a_{ij}\}$ ($N=2^{n}$, $M=2^{m}$), including adding one row to another and swapping two rows.
2.1 Adding one row of a matrix to another row
For the quantum algorithm that adds one row of a matrix to another, the quantum circuit diagram is shown in FIG.1. This algorithm requires two pure states and four auxiliary states, and is divided into eight steps.
Figure 1: Quantum circuit diagram for adding one row of a matrix to another row
The following steps give rise to the algorithm of adding the $(k+1)$th row of a matrix $A$ to the $(l+1)$th row with $10\leq l,k\leq N-$ 1.
Step 1: Prepare the initial state
First, introduce an $n$-qubits register $R_{1}$ and an $m$-qubit register $C_{1}$, which are used to enumerate the rows and columns of matrix $A$, respectively. The entries $a_{ij}$ of the matrix $A$ are then encoded into a pure state as follows,
$|\Psi_{1}\rangle=\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}},\;\;~{}\sum_{ij}|a_{ij}|^{2}=1,$
where $|i\rangle$ is the binary representation of $i$.
Next, introduce an $N$-dimensional auxiliary column vector $Z$ (with the elements of the $k+1$ and the $l+1$ rows being $\frac{1}{\sqrt{2}}$, and all other elements being 0), along with an $n$-qubit register $R_{2}$. The elements of $Z$ are then encoded into pure states as follows,
$|\Psi_{2}\rangle=\frac{1}{\sqrt{2}}(|k\rangle_{R_{2}}+|l\rangle_{R_{2}}).$
The initial state of the entire system is then given by
$|\Phi_{0}\rangle$ $=$ $|\Psi_{1}\rangle\otimes|\Psi_{2}\rangle$
$=$ $\frac{1}{\sqrt{2}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}\Big{)}.$
Step 2: Label the two states of $R_{2}$.
In order to accurately separate the useful and useless information terms in $|\Phi_{0}\rangle$ for subsequent calculations, we need to label the two states of $R_{2}$ in $|\Phi_{0}\rangle$ accordingly.
Introduce a 1-qubit auxiliary $B_{1}$ in state $|0\rangle_{B_{1}}$, and the following control operator,
$W^{(1)}_{R_{2}B_{1}}=P_{R_{2}}\otimes\sigma^{(x)}_{B_{1}}+(I_{R_{2}}-P_{R_{2}})\otimes I_{B_{1}},$
where $P_{R_{2}}=|k\rangle_{R_{2}}{{}_{R_{2}}}\langle k|$ is a projection operator, $\sigma^{(x)}_{B_{1}}$ is the standard Pauli matrix, and $I_{B_{1}}$ is the identity operator acting on $B_{1}$. Applying the operator $W^{(1)}_{R_{2}B_{1}}$ to $|\Phi_{0}\rangle|0\rangle_{B_{1}}$, we get
$|\Phi_{1}\rangle$ $=$ $W^{(1)}_{R_{2}B_{1}}|\Phi_{0}\rangle\otimes|0\rangle_{B_{1}}$
$=$ $\frac{1}{\sqrt{2}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|1\rangle_{B_{1}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|0\rangle_{B_{1}}\Big{)}.$
That is, for the term where the state of $R_{2}$ is $|k\rangle_{R_{2}}$, it is labeled by $|1\rangle_{B_{1}}$, and for the term where the state of $R_{2}$ is $|l\rangle_{R_{2}}$, it is labeled by $|0\rangle_{B_{1}}$.
Since the control operator $W^{(1)}_{R_{2}B_{1}}$ has an $n$-qubit control register, it can be expressed by using $O(n)$ Toffoli gates KShV . Therefore, the circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)=O(\log\,N)$.
Step 3: Separate the useful information from the redundant information.
To precisely extract the element located at the $k+1$th row of the second term in $|\Phi_{1}\rangle$, from the terms in $|\Phi_{1}\rangle$ labeled by $|0\rangle_{B_{1}}$ we only extract the terms where $R_{1}$ is in the state $|k\rangle_{R_{1}}$. To do so, we introduce a 1-qubit auxiliary $B_{2}$ in the state $|0\rangle_{B_{2}}$, and the following control operator,
$W^{(2)}_{R_{1}B_{1}B_{2}}$ $=$ $P_{R_{1}B_{1}}\otimes\sigma^{(x)}_{B_{2}}+(I_{R_{1}B_{1}}-P_{R_{1}B_{1}})\otimes I_{B_{2}},$
where $P_{R_{1}B_{1}}=|k\rangle_{R_{1}}|0\rangle_{B_{1}}\;{{}_{R_{1}}}\langle k|{{}_{B_{1}}}\langle 0|$ is a projection operator.
Applying the operator $W^{(2)}_{R_{1}B_{1}B_{2}}$ to $|\Phi_{1}\rangle|0\rangle_{B_{2}}$, we obtain
$|\Phi_{2}\rangle$ $=$ $W^{(2)}_{R_{1}B_{1}B_{2}}|\Phi_{1}\rangle\otimes|0\rangle_{B_{2}}$
$=$ $\frac{1}{\sqrt{2}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|1\rangle_{B_{1}}|0\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|0\rangle_{B_{1}}|1\rangle_{B_{2}}\Big{)}$
$+|g_{1}\rangle_{R_{1}C_{1}R_{2}}|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$
That is, the terms where the state of $R_{1}$ and $B_{1}$ are $|k\rangle_{R_{1}}|0\rangle_{B_{1}}$ are labeled by $|1\rangle_{B_{2}}$, while the other terms are labeled by $|0\rangle_{B_{2}}$. The states $|1\rangle_{B_{1}}|0\rangle_{B_{2}}$ and $|0\rangle_{B_{1}}|1\rangle_{B_{2}}$ in $|\Phi_{2}\rangle$ correspond to the useful information terms required by the algorithm. All other useless information terms are represented by $|g_{1}\rangle_{R_{1}C_{1}R_{2}}$.
Step 4: Use SWAP gate to exchange the states of $R_{1}$ and $R_{2}$ marked by $|1\rangle_{B_{2}}$.
To add the $(k+1)$th row to the $(l+1)$th row, we apply a SWAP gate operation to the terms in $|\Phi_{2}\rangle$ labeled by $|1\rangle_{B_{2}}$, which swaps the states of $R_{1}$ and $R_{2}$. This operation is given by the following gate,
$W^{(3)}_{R_{1}R_{2}B_{2}}=SWAP_{R_{1},R_{2}}\otimes|1\rangle_{B_{2}}\;{{}_{B_{2}}}\langle 1|+I_{R_{1}R_{2}}\otimes|0\rangle_{B_{2}}\;{{}_{B_{2}}}\langle 0|.$
Applying $W^{(3)}_{R_{1}R_{2}B_{2}}$ to $|\Phi_{2}\rangle$ we obtain
$|\Phi_{3}\rangle$ $=$ $W^{(3)}_{R_{1}R_{2}B_{2}}|\Phi_{2}\rangle$
$=$ $\frac{1}{\sqrt{2}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|1\rangle_{B_{1}}|0\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|0\rangle_{B_{1}}|1\rangle_{B_{2}}\Big{)}$
$+|g_{1}\rangle_{R_{1}C_{1}R_{2}}|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$
Note that the $SWAPs$ in the control operator $W^{(3)}_{R_{1}R_{2}B_{2}}$ have common single control qubit and are related to different pairs of qubits. Therefore, they can be applied simultaneously. Consequently, the depth of this operator is $O(1)$ KShV .
Step 5: Label the useful and redundant information.
To effectively avoid mixing useful and useless information terms in the subsequent computation process, we introduce an auxiliary qubit $B_{3}$ in state $|0\rangle_{B_{3}}$, together with the projection operator $P_{B_{1}B_{2}}=|00\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 00|$. We use $B_{3}$ to label the useful and useless information terms in $|\Phi_{3}\rangle$ accordingly by using the following control operator,
$W^{(4)}_{B_{1}B_{2}B_{3}}=P_{B_{1}B_{2}}\otimes\sigma^{(x)}_{B_{3}}+(I_{B_{1}B_{2}}-P_{B_{1}B_{2}})\otimes I_{B_{3}}.$
Applying $W^{(4)}_{B_{1}B_{2}B_{3}}$ to $|\Phi_{3}\rangle|0\rangle_{B_{3}}$ we obtain
$|\Phi_{4}\rangle$ $=$ $W^{(4)}_{B_{1}B_{2}B_{3}}|\Phi_{3}\rangle|0\rangle_{B_{3}}$
$=$ $\frac{1}{\sqrt{2}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|1\rangle_{B_{1}}|0\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|0\rangle_{B_{1}}|1\rangle_{B_{2}}\Big{)}|0\rangle_{B_{3}}$
$+|g_{1}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}}|1\rangle_{B_{3}}.$
In this way, $|0\rangle_{B_{3}}$ ($|1\rangle_{B_{3}}$) labels the useful (useless) information terms.
The computational depth of the control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ is $O(2)$.
Step 6: Use a Hadamard gate for information superposition.
Applying the Hadamard operator $W^{(5)}_{B_{1}B_{2}}=H^{\otimes 2}$ to $B_{1}$ and $B_{2}$, we obtain
$|\Phi_{5}\rangle$ $=$ $W^{(5)}_{B_{1}B_{2}}|\Phi_{4}\rangle$
$=$ $\frac{1}{(\sqrt{2})^{3}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}|k\rangle_{R_{2}}|0\rangle_{B_{1}}|0\rangle_{B_{2}}|0\rangle_{B_{3}}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}B_{3}}.$
Here, the terms in the state $|0\rangle_{B_{1}}|0\rangle_{B_{2}}|0\rangle_{B_{3}}$ are retained, which contain all the information required for further process. The remaining terms are represented as $|g_{2}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}B_{3}}$.
The computational depth of the control operator $W^{(5)}_{B_{1}B_{2}}$ is $2$.
Step 7: Relabel the useful and redundant information.
Introduce an auxiliary qubit $B_{4}$ in state $|0\rangle_{B_{4}}$, along with the projection operator $P_{B_{1}B_{2}B_{3}}=|000\rangle_{B_{1}B_{2}B_{3}}\;{}_{B_{1}B_{2}B_{3}}\langle 000|$. Applying the control operator
$W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}=P_{B_{1}B_{2}B_{3}}\otimes\sigma^{(x)}_{B_{4}}+(I_{B_{1}B_{2}B_{3}}-P_{B_{1}B_{2}B_{3}})\otimes I_{B_{4}}$
to $|\Phi_{5}\rangle|0\rangle_{B_{4}}$, we obtain
$|\Phi_{6}\rangle$ $=$ $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}|\Phi_{5}\rangle|0\rangle_{B_{4}}$
$=$ $\frac{1}{(\sqrt{2})^{3}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}|k\rangle_{R_{2}}|0\rangle_{B_{1}}|0\rangle_{B_{2}}|0\rangle_{B_{3}}|1\rangle_{B_{4}}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}B_{3}}|0\rangle_{B_{4}}.$
The useful information terms are now labeled by $|1\rangle_{B_{4}}$, while the useless information terms are labeled by $|0\rangle_{B_{4}}$.
The computational depth of the control operator $W^{(6)}_{R_{2}B_{1}B_{2}B_{3}B_{4}}$ is $O(3)$.
Step 8: Measure the auxiliary state $B_{4}$ using $|1\rangle_{B_{4}}$.
By measuring the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$, we get
$|\Phi_{7}\rangle$ $=$ $|\Psi_{out}\rangle|k\rangle_{R_{2}}|000\rangle_{B_{1}B_{2}B_{3}},$ (8)
where
$|\Psi_{out}\rangle=\frac{1}{G}\Big{(}\sum_{i\neq l}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}\\ +\sum_{j=0}^{M-1}(a_{lj}+a_{kj})|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}$
with the normalization
$G=\big{(}\sum_{i\neq l}\sum_{j}|{a_{ij}}|^{2}+\sum_{j}|{a_{kj}+a_{lj}}|^{2}\big{)}^{1/2}.$
The success probability is $\frac{G^{2}}{8}$. This probability clearly depends only on the matrix elements ${a_{ij}}$ and is independent of the matrix dimension.
The overall computational depth of the algorithm is $O(n)$, determined by $W^{(1)}_{R_{2}B_{1}}$ and $W^{(2)}_{R_{1}B_{1}B_{2}}$.
2.2 Swapping two rows of a matrix
For the quantum algorithm that swaps two rows of a matrix, the quantum circuit diagram is shown in FIG.2. The algorithm requires two pure states and four auxiliaries, and is divided into eight steps to swap the $(k+1)$th and $(l+1)$th rows of matrix $A$ ($10\leq l,k\leq N-$ 1).
Figure 2: The quantum circuit diagram for swapping two rows of a matrix.
Step 1: Prepare the initial state.
First, introduce an $n$-qubit register $R_{1}$ and an $m$-qubit register $C_{1}$, which enumerate the rows and columns of the matrix $A$, respectively. The elements of the matrix $A$ are then encoded into the following pure state,
$|\Psi_{1}\rangle=\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}},\;\;\sum_{ij}|a_{ij}|^{2}=1,$
Next, introduce an $N\times N$ auxiliary matrix $Z$ such that the elements in the $(l+1)$-th row and $(k+1)$-th column, the $(k+1)$-th row and $(k+1)$-th column, and the $(l+1)$-th row and $(l+1)$-th column are $\frac{1}{\sqrt{3}}$, the rest elements are 0; along with two $n$-qubit registers $R_{2}$ and $C_{2}$. The elements of $Z$ are then encoded into the following pure state,
$|\Psi_{2}\rangle=\frac{1}{\sqrt{3}}(|l\rangle_{R_{2}}|k\rangle_{C_{2}}+|k\rangle_{R_{2}}|k\rangle_{C_{2}}+|l\rangle_{R_{2}}|l\rangle_{C_{2}}).$
The initial state is of the form,
$|\Phi_{0}\rangle$ $=$ $|\Psi_{1}\rangle\otimes|\Psi_{2}\rangle$
$=$ $\frac{1}{\sqrt{3}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|k\rangle_{C_{2}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|l\rangle_{C_{2}}\Big{)}.$
Step 2: Label the distinct states of $R_{2}$ and $C_{2}$.
To accurately separate the useful from the useless information terms in $|\Phi_{0}\rangle$, introduce an auxiliary qubit $B_{1}$ in state $|0\rangle_{B_{1}}$, and the following control operator,
$W^{(1)}_{R_{2}C_{2}B_{1}}=P_{R_{2}C_{2}}\otimes\sigma^{(x)}_{B_{1}}+(I_{R_{2}C_{2}}-P_{R_{2}C_{2}})\otimes I_{B_{1}},$
where $P_{R_{2}C_{2}}=|l\rangle_{R_{2}}|k\rangle_{C_{2}}{{}_{R_{2}}}\langle l|{{}_{C_{2}}}\langle k|$ is the projection operator. Applying $W^{(1)}_{R_{2}C_{2}B_{1}}$ to $|\Phi_{0}\rangle|0\rangle_{B_{1}}$, we get
$|\Phi_{1}\rangle$ $=$ $W^{(1)}_{R_{2}C_{2}B_{1}}|\Phi_{0}\rangle\otimes|0\rangle_{B_{1}}$
$=$ $\frac{1}{\sqrt{3}}\Big{(}\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|1\rangle_{B_{1}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}$
$+\sum_{i=0}^{N-1}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|l\rangle_{C_{2}}|0\rangle_{B_{1}}\Big{)}.$
In this way, the terms where the states of $R_{2}$ and $C_{2}$ are different are labeled by $|1\rangle_{B_{1}}$, and the terms where the states of $R_{2}$ and $C_{2}$ are the same are labeled by $|0\rangle_{B_{1}}$.
The circuit complexity for computing $|\Phi_{1}\rangle$ is $O(n)=O(\log\,N)$.
Step 3: Separate the useful information terms from the useless information terms.
We proceed to separate the useful information terms from the useless ones.
1) In the first sum, extract all elements except for those in the $(k+1)$th and $(l+1)$th rows.
2) In the second sum, extract all elements in the $(l+1)$th row.
3) In the third sum, extract all elements in the $(k+1)$th row.
4) For the terms in $|\Phi_{1}\rangle$ labeled by $|1\rangle_{B_{1}}$, extract all terms except those where $R_{1}$ is in the states $|k\rangle_{R_{1}}$ and $|l\rangle_{R_{1}}$.
5) For the two sums in $|\Phi_{1}\rangle$ labeled by $|0\rangle_{B_{1}}$, extract the terms where $R_{1}$ is in the state $|l\rangle_{R_{1}}$ and the terms where $R_{1}$ is in the state $|k\rangle_{R_{1}}$.
We introduce an auxiliary qubit $B_{2}$ in state $|0\rangle_{B_{2}}$, and use the following control operators,
$W^{(1)}$ $=$ $P_{R_{1}R_{2}}\otimes\sigma^{(x)}_{{B_{2}}^{1}}+(I_{R_{1}R_{2}}-P_{R_{1}R_{2}})\otimes I_{{B_{2}}^{1}},$
$W^{(2)}$ $=$ $P_{R_{1}C_{2}}\otimes\sigma^{(x)}_{{B_{2}}^{2}}+(I_{R_{1}C_{2}}-P_{R_{1}C_{2}})\otimes I_{{B_{2}}^{1}},$
where $P_{R_{1}R_{2}}=|k\rangle_{R_{1}}|l\rangle_{R_{2}}\;{{}_{R_{1}}}\langle k|{{}_{R_{2}}}\langle l|$ and $P_{R_{1}C_{2}}=|l\rangle_{R_{1}}|k\rangle_{C_{2}}\;{{}_{R_{1}}}\langle l|{{}_{C_{2}}}\langle k|$ are projection operators. $\sigma^{(x)}_{{B_{2}}^{i}}$ and $I_{{B_{2}}^{i}}$ are the operators acting on the $i$-th qubit of the auxiliary $B_{2}$ $(i=1,2)$.
When $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}=W^{(2)}W^{(1)}$ acts on $|\Phi_{1}\rangle|0\rangle_{B_{2}}$, we obtain
$|\Phi_{2}\rangle$ $=$ $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}|\Phi_{1}\rangle\otimes|0\rangle_{B_{2}}$
$=$ $\frac{1}{\sqrt{3}}\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{lj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}|k\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|01\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|l\rangle_{C_{2}}|0\rangle_{B_{1}}|10\rangle_{B_{2}}\Big{)}$
$+|g_{1}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}.$
That is, the terms labeled by $|1\rangle_{B_{1}}|00\rangle_{B_{2}}$, $|0\rangle_{B_{1}}|01\rangle_{B_{2}}$ and $|0\rangle_{B_{1}}|10\rangle_{B_{2}}$ are the useful information terms required by the algorithm, while the other terms are the useless information terms denoted by $|g_{1}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}$.
The computational depth of the control operator $W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}$ is $O(2n)=O(\log\,N)$.
Step 4: Use SWAP gates to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$.
To swap the $(k+1)$th and $(l+1)$th rows of the matrix, we apply SWAP gate to the terms in $|\Phi_{2}\rangle$ labeled by $|10\rangle_{B_{2}}$ and $|01\rangle_{B_{2}}$, so as to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$. Denote
$W^{(1)}$ $=$ $SWAP_{R_{1},C_{2}}\otimes|1\rangle_{{B_{2}}^{1}}\;{{}_{{B_{2}}^{1}}}\langle 1|+I_{R_{1}C_{2}}\otimes|0\rangle_{{B_{2}}^{1}}\;{{}_{{B_{2}}^{1}}}\langle 0|,$
$W^{(2)}$ $=$ $SWAP_{R_{1},R_{2}}\otimes|1\rangle_{{B_{2}}^{2}}\;{{}_{{B_{2}}^{2}}}\langle 1|+I_{R_{1}R_{2}}\otimes|0\rangle_{{B_{2}}^{2}}\;{{}_{{B_{2}}^{2}}}\langle 0|,$
where $|1\rangle_{{B_{2}}^{1}}$ stands for that the first qubit of $B_{2}$ is in the state $|1\rangle$, and $|1\rangle_{{B_{2}}^{2}}$ for that the second qubit of $B_{2}$ in the state $|1\rangle$.
Applying $W^{(3)}_{R_{1}R_{2}C_{2}B_{2}}=W^{(2)}W^{(1)}$ to $|\Phi_{2}\rangle$, we obtain
$|\Phi_{3}\rangle$ $=$ $W^{(3)}_{R_{1}R_{2}C_{2}B_{2}}|\Phi_{2}\rangle$
$=$ $\frac{1}{\sqrt{3}}\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|01\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|10\rangle_{B_{2}}\Big{)}+$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}.$
The SWAP gates transfer the elements originally in the $(k+1)$th row to the $(l+1)$th row in the terms labeled by $|0\rangle_{B_{1}}|01\rangle_{B_{2}}$, and the elements originally in the $(l+1)$th row to the $(k+1)$th row in the terms labeled by $|0\rangle_{B_{1}}|10\rangle_{B_{2}}$.
The circuit complexity for computing $W^{(3)}_{R_{1}R_{2}B_{2}}$ is $O(1)$.
Step 5: Label the useful and useless information terms.
Introduce an auxiliary qubit $B_{3}$ in state $|0\rangle_{B_{3}}$. To use $B_{3}$ to label the useful and useless information terms in $|\Phi_{3}\rangle$, we apply the following control operator,
$W^{(4)}_{B_{1}B_{2}B_{3}}=V^{(1)}V^{(2)}V^{(3)},$
where $V^{(i)}=P^{(i)}_{B_{1}B_{2}}\otimes\sigma^{(x)}_{B_{3}}+(I_{B_{1}B_{2}}-P_{B_{1}B_{2}})\otimes I_{B_{3}}$, with the projection operators $P^{(i)}_{B_{1}B_{2}}$ $(i=1,2,3)$ given by
$P^{(1)}_{B_{1}B_{2}}$ $=$ $|100\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 100|,$
$P^{(2)}_{B_{1}B_{2}}$ $=$ $|001\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 001|,$
$P^{(3)}_{B_{1}B_{2}}$ $=$ $|010\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 010|.$
Acting $W^{(4)}_{B_{1}B_{2}B_{3}}$ on $|\Phi_{3}\rangle|0\rangle_{B_{3}}$, we obtain
$|\Phi_{4}\rangle$ $=$ $W^{(4)}_{B_{1}B_{2}B_{3}}|\Phi_{3}\rangle|0\rangle_{B_{3}}$
$=$ $\frac{1}{\sqrt{3}}\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|01\rangle_{B_{2}}$
$+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|10\rangle_{B_{2}}\Big{)}|1\rangle_{B_{3}}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}|0\rangle_{B_{3}}.$
Here, $|1\rangle_{B_{3}}$ labels the useful information term, while $|0\rangle_{B_{3}}$ labels the useless information term.
The control operator $W^{(4)}_{B_{1}B_{2}B_{3}}$ has a computational depth of $3$.
Step 6: Use Hadamard gates for information superposition.
Applying the Hadamard operator $W^{(5)}_{B_{1}B_{2}}=H^{\otimes 3}$ to $B_{1}$ and $B_{2}$, we have
$|\Phi_{5}\rangle$ $=$ $W^{(5)}_{B_{1}B_{2}}|\Phi_{4}\rangle$
$=$ $\frac{1}{(\sqrt{2})^{3}\cdot\sqrt{3}}\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}$
$|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0\rangle_{B_{1}}|00\rangle_{B_{2}}|1\rangle_{B_{3}}+|g_{3}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}B_{3}},$
where the terms in the state $|0\rangle_{B_{1}}|00\rangle_{B_{2}}|1\rangle_{B_{3}}$ are retained, the other terms are denoted by $|g_{3}\rangle_{R_{1}C_{1}R_{2}B_{1}B_{2}B_{3}}$.
At this point, we have successfully swaped the elements of the $(k+1)$-th row of matrix $A$ with the corresponding elements of the $(l+1)$-th row. The control operator $W^{(5)}_{B_{1}B_{2}}$ has a computational depth of $3$.
Step 7: Re-label the useful information terms and useless information terms.
Since new useless information terms are generated in the previous operation, and the useful information is stored in the first term of $|\Phi_{5}\rangle$, we need to re-mark the useful and useless information terms. To do so, we introduce an auxiliary qubit $B_{4}$ in state $|0\rangle_{B_{4}}$ and a projection operator $P_{B_{1}B_{2}B_{3}}=|0001\rangle_{B_{1}B_{2}B_{3}}\;{{}_{B_{1}B_{2}B_{3}}}\langle 0001|.$ We apply the following control operator
$W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}=P_{B_{1}B_{2}B_{3}}\otimes\sigma^{(x)}_{B_{4}}+(I_{B_{1}B_{2}B_{3}}-P_{B_{1}B_{2}B_{3}})\otimes I_{B_{4}}$
to state $|\Phi_{5}\rangle|0\rangle_{B_{4}}$, which gives rise to
$|\Phi_{6}\rangle$ $=$ $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}|\Phi_{5}\rangle|0\rangle_{B_{4}}$
$=$ $\frac{1}{(\sqrt{2})^{3}\cdot\sqrt{3}}\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}|l\rangle_{R_{2}}|k\rangle_{C_{2}}$
$|0\rangle_{B_{1}}|00\rangle_{B_{2}}|1\rangle_{B_{3}}|1\rangle_{B_{4}}+|g_{3}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}B_{3}}|0\rangle_{B_{4}},$
where the useful information terms are marked by $|1\rangle_{B_{4}}$, and the useless information terms are marked by $|0\rangle_{B_{4}}$.
The control operator $W^{(6)}_{B_{1}B_{2}B_{3}B_{4}}$ has a computational depth of $4$.
Step 8: Measure the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$
We measure the auxiliary $B_{4}$ in the state $|1\rangle_{B_{4}}$ to remove the useless information terms, and obtain
$|\Phi_{7}\rangle=|\Psi_{out}\rangle|l\rangle_{R_{2}}|k\rangle_{C_{2}}|0001\rangle_{B_{1}B_{2}B_{3}},$ (16)
where
$|\Psi_{out}\rangle$ $=$ $\Big{(}\sum_{i\neq l,k}\sum_{j=0}^{M-1}a_{ij}|i\rangle_{R_{1}}|j\rangle_{C_{1}}$
$+\sum_{j=0}^{M-1}a_{lj}|k\rangle_{R_{1}}|j\rangle_{C_{1}}+\sum_{j=0}^{M-1}a_{kj}|l\rangle_{R_{1}}|j\rangle_{C_{1}}\Big{)}.$
Clearly, the success probability of the measurement is $\frac{1}{24}$ and the overall depth of the algorithm is $O(n)$.
3 Matrix Trace Calculation and Matrix Transposition
In this section we present quantum algorithms for trace calculation of an $N\times N$ matrix and transposition of an $N\times M$ matrix.
3.1 Matrix Trace Calculation
The quantum circuit for calculating the trace of the matrix $S$ is shown in FIG.3. Assume that a matrix $S$=$\{a_{ik}\}$ is of size $N\times N$ with $N=2^{n}$. This algorithm requires 1 pure state and 3 auxiliary qubits, and is divided into the following 6 steps.
Figure 3: Quantum circuit for calculating matrix trace.
Step 1: Prepare the initial state.
First, introduce two $n$-qubit registers $R$ and $C$, which are used to enumerate the rows and columns of the matrix $S$, and encode the elements of the matrix $S$ to obtain the following pure initial state,
$|\Phi_{0}\rangle=\sum_{i=0}^{N-1}\sum_{k=0}^{N-1}a_{ik}|i\rangle_{R}|k\rangle_{C},\;\;{\sum_{ik}|a_{ik}|^{2}=1}.$ (18)
Step 2: Mark the terms where the states of $R$ and $C$ are the same.
To calculate the sum of the diagonal elements of matrix $S$, we need to mark the diagonal elements of $S$, i.e., the terms where the states of registers $R$ and $C$ are the same. To do this, introduce an $n$-qubit auxiliary $A$ in the state $|0\rangle_{A}$, and an operator
$W_{j}^{(m)}$ $=$ $P^{(m)}_{j}\otimes\sigma^{(x)}_{A^{(j)}}+(I_{j}-P^{(m)}_{j})\otimes I_{A^{(j)}},$
where $P^{(m)}_{j}=|m_{j}\rangle_{R}|m_{j}\rangle_{C}\;{{}_{R}}\langle m_{j}|{{}_{C}}\langle m_{j}|$, $m=0,1$, $j=1,2,...,n$, is the projection operator acting on the $j$-th qubit of registers $R$ and $C$, $\sigma^{(x)}_{A^{(j)}}$ is the Pauli matrix acting on the $j$-th qubit of auxiliary $A$, $I_{A^{(j)}}$ is the identity operator acting on the $j$-th qubit of auxiliary $A$.
Applying the operator $W_{RCA}^{(1)}=\prod_{j=1}^{n}W_{j}^{0}W_{j}^{1}$ to $|\Phi_{0}\rangle|0\rangle_{A}$, we obtain
$|\Phi_{1}\rangle$ $=$ $W^{(1)}_{RCA}|\Phi_{0}\rangle|0\rangle_{A}$
$=$ $\left(\sum_{i=0}^{N-1}a_{ii}|i\rangle_{R}|i\rangle_{C}\right)|N-1\rangle_{A}+|g_{1}\rangle_{RCA}.$
In the process of realizing the quantum state $|\Phi_{1}\rangle$, we check whether the $j$-th qubits of registers $R$ and $C$ are the same. If their $j$-th qubits are the same, the $j$-th qubit of $A$ is flipped from $|0\rangle$ to $|1\rangle$. If the states of all $n$ qubits of $R$ and $C$ are the same, the states of all $n$ qubits of $A$ are flipped from $|0\rangle_{A}$ to $|N-1\rangle_{A}$.
Step 3: Mark the useful and useless information terms.
Currently, all useful information terms are stored in the first sum of $|\Phi_{1}\rangle$, while the second sum $|g_{1}\rangle_{{RCA}}$ represents useless information terms that are not needed by the algorithm.
To separately mark the useful and useless information terms in the state $|\Phi_{1}\rangle$, we introduce the projection operator $P_{A}=|N-1\rangle_{A}\;{{}_{A}}\langle N-1|,$ and a 1-qubit auxiliary $B_{1}$ in the state $|0\rangle_{B_{1}}$. The control operator $W^{(2)}_{AB_{1}}=P_{A}\otimes\sigma^{(x)}_{B_{1}}+(I_{A}-P_{A})\otimes I_{B_{1}}$ is then applied to $A$ and $B_{1}$, giving rise to
$|\Phi_{2}\rangle$ $=$ $W^{(2)}_{AB_{1}}|\Phi_{1}\rangle|0\rangle_{B_{1}}$
$=$ $\left(\sum_{i=0}^{N-1}a_{ii}|i\rangle_{R}|i\rangle_{C}\right)\otimes|N-1\rangle_{A}\otimes|1\rangle_{B_{1}}$
$+|g_{1}\rangle_{RCA}\otimes|0\rangle_{B_{1}}.$
Here, the useful information terms are marked by $|1\rangle_{B_{1}}$, and the useless information terms are marked by $|0\rangle_{B_{1}}$.
The computational depth for obtaining $|\Phi_{2}\rangle$ is $O(n)=O(\log\,N)$.
Step 4: Use the Hadamard gate for information superposition.
Up to this point, we have obtained the diagonal elements of $S$. To compute the sum we apply the Hadamard operator $W^{(3)}_{RCA}=H^{\otimes 3n}$ to all qubits in $|\Phi_{2}\rangle$ except for the auxiliary $B_{1}$. We get
$|\Phi_{3}\rangle$ $=$ $W^{(3)}_{RCA}|\Phi_{2}\rangle$
$=$ $\frac{1}{2^{3n/2}}\left(\sum_{i=0}^{N-1}a_{ii}\right)|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}+|g_{2}\rangle_{RCAB_{1}}.$
In the state $|\Phi_{3}\rangle$, we select the terms in the state $|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}$, as these terms contain all the useful information required for the trace calculation. The other terms are represented by $|g_{2}\rangle_{RCAB_{1}}$.
Step 5: Re-mark the useful and useless information terms.
Since new useless information terms were generated in the previous step and the useful information is stored in the first term of $|\Phi_{3}\rangle$, we need to re-mark the useful and useless information terms. To do this, we introduce an auxiliary qubit $B_{2}$ in state $|0\rangle_{B_{2}}$, as well as the projection operator $P_{RCAB_{1}}=|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}\;{{}_{R}}\langle 0|{{}_{C}}\langle 0|_{A}\langle 0|_{B_{1}}\langle 1|.$ We apply the following control operator
$W^{(4)}_{RCAB_{1}B_{2}}=P_{RCAB_{1}}\otimes\sigma^{(x)}_{B_{2}}+(I_{RCAB_{1}}-P_{RCAB_{1}})\otimes I_{B_{2}}\,$
to $|\Phi_{3}\rangle\otimes|0\rangle_{B_{2}}$. We obtain
$|\Phi_{4}\rangle$ $=$ $W^{(4)}_{RCAB_{1}B_{2}}|\Phi_{3}\rangle|0\rangle_{B_{2}}$
$=$ $\frac{1}{2^{3n/2}}\left(\sum_{i=0}^{N-1}a_{ii}\right)|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}|1\rangle_{B_{2}}$
$+|g_{2}\rangle_{RCAB_{1}}|0\rangle_{B_{2}}.$
Here, the useful information terms are marked by $|1\rangle_{B_{2}}$, and the useless information terms are marked by $|0\rangle_{B_{2}}$.
The control operator $W^{(4)}_{S_{1}S_{2}AB_{1}B_{2}}$ has $13n+$ 1 control qubits, and can be represented using $O(3n)=O(n)$ Toffoli gates. So it takes $O(n)=O(\log\,N)$ steps to compute $|\Phi_{5}\rangle$.
Step 6: Measure the auxiliary qubit $B_{2}$ in the state $|1\rangle_{B_{2}}$.
The trace of the matrix $S$ is stored in the quantum state $|\Phi_{4}\rangle$ in the form of probability amplitudes, with the useful information marked by $|1\rangle_{B_{2}}$. To extract the useful information, we measure the auxiliary qubit $B_{2}$ in the state $|1\rangle_{B_{2}}$, and obtain
$|\Phi_{5}\rangle$ $=$ $\frac{\sum_{i=0}^{N-1}a_{ii}}{|\sum_{i=0}^{N-1}a_{ii}|}|0\rangle_{R}|0\rangle_{C}|0\rangle_{A}|1\rangle_{B_{1}}.$ (23)
The trace is obtained based on the measurement probability $\frac{|\sum_{i=0}^{N-1}a_{ii}\rangle|^{2}}{2^{3n}}$.
The overall depth of the algorithm given by the operators $W^{(2)}_{AB_{1}}$ and $W^{(4)}_{RCAB_{1}B_{2}}$ is $O(n)=O(logN)$.
3.2 Matrix Transposition
The quantum circuit for matrix transposition is shown in FIG.4. Assume that a matrix $S$=$\{a_{ik}\}$ is of size $N\times M$ with $N=2^{n}$ and $M=2^{m}$. The algorithm requires 2 pure states and 1 auxiliary qubit, and is divided into 3 steps.
Figure 4: Quantum Circuit for Matrix Transposition.
Step 1: Prepare the initial state.
First, introduce an $n$-qubit register $R$ and an $m$-qubit register $C$, where $R$ and $C$ enumerate the rows and columns of the matrix $S$, respectively. The elements of matrix $S$ are encoded into the following pure state,
$|\Psi\rangle=\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|i\rangle_{R}|k\rangle_{C},\;\;\sum_{ik}|a_{ik}|^{2}=1,$
Then, introduce a pure state $|0\rangle_{D}$ for an $m$-qubit register $D$, and an auxiliary qubit $B$ in state $|0\rangle_{B}$. The initial state is give by
$|\Phi_{0}\rangle$ $=$ $|0\rangle_{D}\otimes|\Psi\rangle\otimes|0\rangle_{B}$
$=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|0\rangle_{D}|i\rangle_{R}|k\rangle_{C}|0\rangle_{B}.$
Step 2: Use a SWAP gate to exchange the states of $D$ and $C$.
We use a SWAP gate to exchange the states of the $D$ and $C$ registers. The operator $W_{DC}=SWAP_{DC}\otimes|0\rangle_{B}\;{{}_{B}}\langle 0|+I_{DC}\otimes|1\rangle_{B}\;{{}_{B}}\langle 1|$ is applied to $|\Phi_{0}\rangle$, yielding
$|\Phi_{1}\rangle$ $=$ $W_{DC}|\Phi_{0}\rangle$
$=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|i\rangle_{D}|k\rangle_{R}|0\rangle_{C}|0\rangle_{B}.$
Step 3: Measure the auxiliary qubit $B$ in the state $|0\rangle_{B}$.
By measuring the auxiliary qubit $B$ in the state $|0\rangle_{B}$, we extract the information of the transposed matrix $S$. This gives the final state:
$|\Phi_{2}\rangle$ $=$ $|\Psi_{out}\rangle|0\rangle_{C},$ (26)
where
$|\Psi_{out}\rangle$ $=$ $\Big{(}\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|i\rangle_{D}|k\rangle_{R}\Big{)}.$ (27)
Clearly, the success probability of this measurement is 1.
By the analysis above, we conclude that the overall depth of the algorithm is $O(1)$, which determined by the operator $W_{DC}$.
4 Conclusion
We have developed quantum algorithms for two types of elementary row transformations of matrices, the trace and transpose of matrices. The operational mechanism of all these algorithms is as follows: First, construct the initial state of the entire system. Second, utilize unitary operations and ancillary qubits to mark the information items required by the algorithm. Third, apply the Hadamard transformation to superpose these information items. Finally, measure the ancillary qubits to eliminate the useless information items and retain only the useful ones.
In all these algorithms, the final result is stored in the probability amplitude of certain quantum states. Our success probability for adding one row of a matrix to another is $\frac{G^{2}}{8}$. The success probability for swapping two rows of a matrix is $\frac{1}{24}$, while the success probability of matrix transposition is 1. The success probability for the trace calculation is $\frac{|\langle\sum_{i=0}^{N-1}a_{ii}\rangle|^{2}}{2^{3n}}$. The success probabilities either depend solely on the elements of the matrix itself, or are fixed or related to the dimensions of the matrix. The larger the dimension, the smaller the success probability. To address this challenge, amplitude amplification techniques may be employed. It is worth noting that the computational depth of these algorithms is related to the dimension of the matrices. It increases logarithmically as the matrix dimension grows. Specifically, the computational depth is $O(n)$ for elementary row operations and trace calculation, and $O(1)$ for matrix transposition.
It would be desired to develop new methods to improve the success probability given by the final measurement in the algorithm, or explore the designs of other quantum algorithms for matrix operations based on multi-qubit Toffoli type gates and the simplest single qubit operations. It would be also interesting to conduct further researches on quantum algorithms for other matrix operations.
Acknowledgments
This work is being supported by the following: the National Natural Science Foundation of China (NSFC) under Grants 11761073, 12075159, and 12171044; the Academician Innovation Platform of Hainan Province. We are deeply grateful to Prof. Junde Wu for his active participation and invaluable contributions to this research.